OpenStudy (anonymous):

If x and y are prime numbers where unity equals the the square of the prime number "x" minus the twice of the square of the prime number y , then find the prime numbers x and y

OpenStudy (anonymous):

@Kainui

OpenStudy (anonymous):

this was the question i was working on

OpenStudy (kainui):

x^2-2y^2=1 Ohhhh I see. This is similar to that other one you were working on a week or two ago that one was great fun but it was p^2-q^2=r and what are all the prime numbers that satisfy this equation I believe, right?

OpenStudy (anonymous):

I have a bad memory lol , i don't remember

OpenStudy (kainui):

I was gonna say if it was simply x^2-y^2=1 then we can pretty quickly guess x=3 and y=2 haha.

OpenStudy (anonymous):

Yeah hehe

OpenStudy (kainui):

y must be 2 though.

OpenStudy (kainui):

for this question as well. Interesting.

OpenStudy (anonymous):

then x = 3

OpenStudy (rational):

x^2 -1 = 2y^2 (x-1)(x+1) = 2y^2 clearly x needs to have an odd parity and (3, 2) is one parituclar solution For a general method for finding all the solutions of any equation of form `x^2-ny^2=1`, you need to study Pell's equation : http://en.wikipedia.org/wiki/Pell's_equation

OpenStudy (anonymous):

I did it without pell's equation

OpenStudy (kainui):

wait I had guessed the right answer I had been thinking 3^2-2^3=1

OpenStudy (rational):

how exactly you did it and are you sure (3,2) is the only possible solution ?

OpenStudy (anonymous):

Trial and error works for small powers and small number though , but good in MCQ's

OpenStudy (anonymous):

@rational yes

OpenStudy (rational):

good, make me feel sure too :) cuz i don't see yet how (3,2) is the only possible solution :/

OpenStudy (anonymous):